# Characterizing spin ordering via maximal row correlation in classical spin models

### v1

@@ III.3 $q=2$ and $q=3$ AFM Potts model on Union-Jack lattice
Figure 6: (a) The Union-Jack lattice has a tripartite structure, formed by inserting additional sites (red circles) at the centers of the plaquettes of a square lattice (blue circles). Each red site connects to its four nearest blue neighbors. The three sublattices have coordinate numbers $8$, $8$, and $4$, respectively, resulting in an average coordinate number $z=6$. To map the spin configuration on the lattice into matrix form, an orthogonal coordinate frame is employed. A snapshot of the spin configuration on $L=16$ Union-Jack lattice for the $q=2$ AFM Potts model is shown in (b), with two colors (red and blue) presenting two spin values. It is taken from the MC sampling after the equilibration at the temperature $T=1.0\times 10^{-5}$. We observe that due to the frustration, the absence of discernible order in the low-temperature configuration makes it challenging to identify an appropriate order parameter.
@@ IV Summary and discussion
In this study, we demonstrate that the quantity of the maximal row correlation $O$, whose magnitude is proportional to the length of the longest vectors obtainable by linear superposition of the row vectors in the spin configuration matrix, captures a key aspect of the intrinsic spin correlations within a given configuration. MC simulations on the Potts models confirm the effectiveness of $O$ as an order parameter. Specifically, we find that $O$ approaches a finite value as $T\to 0$ whenever some degree of spin ordering (either partial or full) is present, regardless of the underlying lattice geometry, interaction type (FM or AFM), or the presence of frustration. In the high-temperature limit where spin order is destroyed, $O$ vanishes consistently. The onset of this critical behavior in $O$—or, more sharply, the peak position of its fluctuations $\delta O$—can be used to accurately identify the critical temperature of the phase transition. These results establish a general and robust framework for investigating phase transitions in classical spin systems, with $O$ serving as a versatile and broadly applicable order parameter.
@@
We further show that the critical behavior of $O$, especially the associated critical exponents, closely parallels that of conventional magnetization—whether full, partial, or staggered—which typically requires careful definition in partially ordered or frustrated systems. As anticipated, the fluctuation $\delta O$ appears to correspond to a susceptibility-like quantity. This correspondence has been quantitatively confirmed in the FM cases (results not shown). However, accurately extracting the associated critical exponents in systems with irregular lattices or frustration requires more extensive simulations on larger system sizes. We defer this investigation to future work.
@@
We conclude this section by speculating potential applications of the order parameter $O$ introduced in this study. As discussed, $O$ is expected to serve as a unified order parameter applicable to a broad class of spin systems. Notably, since $O$ is derived from the row correlation matrix—specifically, it is proportional to the square root of the matrix’s largest eigenvalue—its inherently nonlocal character may offer particular utility in identifying exotic phases that lack local order but exhibit nontrivial global structures. Moreover, because our approach operates directly on configurations generated via MC sampling and does not rely on prior knowledge of the system’s symmetry or structure, it is naturally well-suited for integration with machine learning techniques [18, 27, 28]. Their combination presents a promising direction for future research.
@@ Acknowledgment
We thank Liping Yang for helpful discussions. Y.-Y.T. and H.L. acknowledge support from the National Natural Science Foundation of China (NSFC; Grants No. 12174168 and No. 12247101).
@@ A.2.1 Full long-range order
A.2.2 Partial order in the $q=3$ AFM Potts model on diced lattice
In the low-temperature limit, partial order in the $q=3$ AFM Potts model on diced lattice emerges on the minority sublattice (blue circles in Fig. 4(a)), where spins adopt a fixed value. In contrast, spins on the majority sublattice (red circles forming a honeycomb sublattice) fluctuate between the remaining two values. As a result, on average the row correlation matrix $M$ under the oblique rectilinear coordinate frame (see Fig. 4(b)) takes the form:
@@ A.2.2 Partial order in the $q=3$ AFM Potts model on diced lattice
The expectation values of the matrix element $m_{ij}$ are easy to calculate since it is simply the inner product of the $i$-th and $j$-th rows of the spin configuration matrix $A$. To calculate the inner product, we need to specify the product between two spin variables, which is given by Eq. (2) for the Potts model. For the diced lattice, when the row indices satisfy $\left|i-j\right|\equiv 1,2\pmod{3}$, a straightforward counting shows that among the $L$ spin pair products appearing in the inner product expression (see Eq. (1)), two-thirds involve a product between an ordered site (with a fixed spin) and a random site (with a random spin), each contributing an expectation value of $-1/2$. The remaining one-third are products between two random sites, yielding an expectation value $\frac{1}{4}(2\times 1+2\times(-\frac{1}{2}))=1/4$. Thus in this situation we have
@@
Appendix B Higher dimensional systems
In this appendix, we apply our method to higher-dimensional cases with $D>2$. For simplicity and illustrative purposes, we consider only the FM Potts model with $q=2$—equivalent to the FM Ising model—on simple cubic lattices in three and four dimensions.
It is important to note that the definition of $O$ requires a slight modification in three dimensions as $O=\sqrt[3]{\lambda_{\mathrm{M}}}/L$, i.e., the cubic root of the largest eigenvalue of $M$, divided by the matrix size $L$. This adjustment is motivated by the fact that, for a fully ordered configuration, the elements of $M$ scale as $L^{2}$, in contrast to the $L$ scaling in the 2D case (cf. Eq. (6)). An analogous procedure applies in 4D, where the last three indices are combined into one, and the order parameter is correspondingly defined as $4]{\lambda_{\mathrm{M}}}/L$.
@@ Appendix B Higher dimensional systems
Our MC results for the $q=2$ FM Potts model on 3D and 4D cubic lattices are shown in Figs. 10 and 11, where the temperature dependences of $O$ and its fluctuation $\delta O$ are plotted for system size up to $L=16$. We see that in both cases, $O$ converges uniformly to unity as $T\to 0$, indicating a fully ordered state at zero temperature. As $L$ increases, the peak in $\delta O$ sharpens and approaches the critical temperature $T_{c}$ previously determined by large-scale MC simulations [29, 30]. Given the limited system sizes accessible in this study—and because a precise determination of critical properties of these models is beyond our present scope—we did not attempt a data collapse to extract $T_{c}$ or the critical exponents $\nu$ and $\beta$.
Appendix C Continuous spin models
In this appendix, we study two paradigmatic models with continuous spin variables on the square lattice: the celebrated XY model and the classical Heisenberg model. The XY model undergoes a Berezinskii-Kosterlitz-Thouless (BKT) phase transition [31, *Kosterlitz_1973, *Kosterlitz_1974] at a critical temperature $T_{c}\approx 0.894$ [34], separating a low-temperature phase with quasi-long-range order from a high-temperature disordered phase characterized by an exponential decay of correlations. In contrast, for the 2D classical Heisenberg model, the existence of a BKT-like transition at finite temperature has been suggested but remains a subject of ongoing debate [35, 36, 37].
@@ Appendix C Continuous spin models
Figures 12 and 13 present preliminary MC results for the XY and classical Heisenberg model on the square lattice, where we compute the temperature dependence of the order parameter $O$ and its fluctuations $\delta O$ for system size up to $L=64$. A notable feature in the behavior of $O$ is observed: unlike previous cases involving systems with fully or partial orders at low temperatures—where $O$ gradually saturates to a finite value and exhibits a clear plateau—here, although $O$ approaches to unity as $T\to 0$, it does so with a significant slope. We interpret this as an indication of the absence of long-range order at any nonzero temperature, consistent with the celebrated Mermin–Wagner–Hohenberg theorem [38, *Hohenberg_1967]. Moreover, the XY and Heisenberg models exhibit remarkable similar behavior. We note that our MC results for $O$ and $\delta O$ closely resemble those previously reported for the magnetization $M$ and susceptibility $\chi$ [35]. It would be worthwhile to pursue more systematic studies at larger system sizes in the future to further substantiate these observations.

### v2

@@ III.3 $q=2$ and $q=3$ AFM Potts model on Union-Jack lattice
Figure 6: (a) The Union-Jack lattice has a tripartite structure, formed by inserting additional sites (red circles) at the centers of the plaquettes of a square lattice (blue circles). Each red site connects to its four nearest blue neighbors. The three sublattices have coordinate numbers 88, 88, and $4$, respectively, resulting in an average coordinate number $z=6$. To map the spin configuration on the lattice into matrix form, an orthogonal coordinate frame is employed. A snapshot of the spin configuration on $L=16$ Union-Jack lattice for the $q=2$ AFM Potts model is shown in (b), with two colors (red and blue) presenting two spin values. It is taken from the MC sampling after the equilibration at the temperature $T=1.0\times 10^{-5}$. We observe that due to the frustration, the absence of discernible order in the low-temperature configuration makes it challenging to identify an appropriate order parameter.
@@ IV Summary and discussion
In this study, we demonstrate that the quantity of the maximal row correlation $O$, whose magnitude is proportional to the length of the longest vectors obtainable by linear superposition of the row vectors in the spin configuration matrix, captures a key aspect of the intrinsic spin correlations within a given configuration. MC simulations on the Potts models confirm the effectiveness of $O$ as an order parameter. Specifically, we find that $O$ approaches a finite value as $T\to 0$ whenever some degree of spin ordering (either partial or full) is present, regardless of the underlying lattice geometry, interaction type (FM or AFM), or the presence of frustration. In the high-temperature limit where spin order is destroyed, $O$ vanishes consistently. The onset of this critical behavior in OO—or, more sharply, the peak position of its fluctuations $\delta O$—can be used to accurately identify the critical temperature of the phase transition. These results establish a general and robust framework for investigating phase transitions in classical spin systems, with $O$ serving as a versatile and broadly applicable order parameter.
@@
We further show that the critical behavior of $O$, especially the associated critical exponents, closely parallels that of conventional magnetization—whether full, partial, or staggered—which typically requires careful definition in partially ordered or frustrated systems. As anticipated, the fluctuation $\delta O$ appears to correspond to a susceptibility-like quantity. This correspondence is quantitatively confirmed for the FM cases and the AFM case on the diced lattice [see Appendix B]. However, accurately extracting the associated critical exponents in systems with irregular lattices or frustration requires more extensive simulations on larger system sizes. We defer this investigation to future work.
@@
We conclude this section by speculating potential applications of the order parameter $O$ introduced in this study. As discussed, $O$ is expected to serve as a unified order parameter applicable to a broad class of spin systems. Notably, since $O$ is derived from the row correlation matrix—specifically, it is proportional to the square root of the matrix’s largest eigenvalue—its inherently nonlocal character may offer particular utility in identifying exotic phases that lack local order but exhibit nontrivial global structures.
Moreover, we note that this technique bears resemblance to PCA, a widely used feature extraction technique in machine learning that has been applied to detect and characterize phase transitions in various models [22, 29, 30, 16, 31, 32]. However, as emphasized in Sec. II.1, the row correlation matrix $M$ differs from the covariance matrices employed in PCA, since it is constructed directly from individual spin configurations. As a result, the order parameter $O$ is closely tied to spin correlations and magnetic orders, offering clear physical interpretability. This framework also enables the extraction of the associated critical exponents through statistical and scaling analysis of $O$ and its fluctuation $\delta O$. Because our approach operates directly on MC-generating configurations without requiring prior knowledge of the system’s symmetry or structure, it may offer novel insights into the mechanisms underlying machine learning–based phase classification [19, 33, 34]. Integrating this method with modern machine learning techniques thus presents a promising avenue for future research.
@@ Acknowledgment
We thank Li-Ping Yang for helpful discussions. We acknowledge support from the National Natural Science Foundation of China (NSFC; Grants No. 12174168 and No. 12247101).
Data Availability
The data that support the findings of this article are openly available [35].
@@ A.2.1 Full long-range order
A.2.2 Partial order in the $q=3$ AFM Potts model on the diced lattice
In the low-temperature limit, partial order in the $q=3$ AFM Potts model on the diced lattice emerges on the minority sublattice (blue circles in Fig. 4(a)), where spins adopt a fixed value. In contrast, spins on the majority sublattice (red circles forming a honeycomb sublattice) fluctuate between the remaining two values. As a result, on average the row correlation matrix $M$ under the oblique rectilinear coordinate frame (see Fig. 4(b)) takes the form:
@@ A.2.2 Partial order in the $q=3$ AFM Potts model on the diced lattice
The expectation values of the matrix element $m_{ij}$ are easy to calculate since it is simply the inner product of the $i$-th and $j$-th rows of the spin configuration matrix $A$. To calculate the inner product, we need to specify the product between two spin variables, which is given by Eq. (2) for the Potts model. For the diced lattice, when the row indices satisfy $\left|i-j\right|\equiv 1,2\pmod{3}$, a straightforward counting shows that among the $L$ spin pair products appearing in the inner product expression (see Eq. (1)), two-thirds involve a product between an ordered site (with a fixed spin) and a random site (with a random spin), each contributing an expectation value of 1/2-1/2. The remaining one-third are products between two random sites, yielding an expectation value $\frac{1}{4}(2\times 1+2\times(-\frac{1}{2}))=1/4$. Thus in this situation we have
@@
Appendix B Data collapse results for $\delta O$
From the previous discussion in Sec. III, we have learned that the critical behavior of the maximal row correlation $O$ is governed by the critical exponent $\beta$, traditionally associated with the magnetization $M$. In this appendix, we further show that, analogously, the scaling behavior of its fluctuations $\delta O$ near the critical temperature can be used to extract the standard critical exponent $\gamma$. As a demonstration, Fig. 10 presents the results of the data collapse of $\delta O$ for the $q=2$ and $q=3$ FM Potts models on the square lattice, and the $q=3$ AFM Potts model on the diced lattice.
Figure 10: Results of the data collapse for the fluctuations of the maximal row correlation $\delta O$ across different models: (a) the $q=2$ FM Potts model on the square lattice; (b) the $q=3$ FM Potts model on the square lattice; (c) the $q=3$ AFM Potts model on the diced lattice.
As shown in Fig. 10, we see that the data collapse for FM interactions is highly satisfactory for both $q=2$ (Ising) and $q=3$ cases. For $q=2$, the collapse aligns well with the exact critical exponents $\nu=1$, and $\gamma=7/4=1.75$ [6] [Fig. 10(a)]. For $q=3$, the known exact values are $\nu=5/6\approx 0.833$ and $\gamma=13/9\approx 1.444$ [5, 6]; our numerical results yield $\nu=0.85(2)$, $\gamma=1.49(2)$ [Fig. 10(b), which are in good agreement within numerical uncertainty.
In contrast, for the $q=3$ Potts model with AFM interactions on the diced lattice, the data collapse is notably less precise when using the same set of system sizes as in the $q=2$ and $q=3$ FM Potts models. The optimal values of $T_{c}$, $\nu$, and $\gamma$ yield a visible more scattered distribution of data points [Fig. 10(c)], resulting in a comparatively large numerical uncertainties associated with the estimates of the critical exponents $\nu$ and $\gamma$. This reduced convergence may be attributed to the geometric irregularity of the diced lattice and the presence of partial ordering fluctuations. Achieving comparable accuracy would likely require simulations on larger system sizes. Nevertheless, the estimated values of $\nu$ and $\gamma$ remain consistent with the conclusion of Ref. [12], which suggests that the universality class of the $q=3$ AFM Potts model on the diced lattice is equivalent to that of the $q=3$ Potts ferromagnet.
Appendix C Higher dimensional systems
@@
In this appendix, we apply our method to higher-dimensional cases with $D>2$. For simplicity and illustrative purposes, we consider only the FM Potts model with q=2q=2—equivalent to the FM Ising model—on simple cubic lattices in three and four dimensions.
It is important to note that the definition of $O$ requires a slight modification in three dimensions as $O=\sqrt[3]{\lambda_{\mathrm{M}}}/L$, i.e., the cubic root of the largest eigenvalue of $M$, divided by the matrix size $L$. This adjustment is motivated by the fact that, for a fully ordered configuration, the elements of $M$ scale as $L^{2}$, in contrast to the $L$ scaling in the 2D case (cf. Eq. (7)). An analogous procedure applies in 4D, where the last three indices are combined into one, and the order parameter is correspondingly defined as $\sqrt[4]{\lambda_{\mathrm{M}}}/L$.
Our MC results for the $q=2$ FM Potts model on 3D and 4D cubic lattices are shown in Figs. 11 and 12, where the temperature dependences of $O$ and its fluctuation $\delta O$ are plotted for system size up to $L=16$. We see that in both cases, $O$ converges uniformly to unity as $T\to 0$, indicating a fully ordered state at zero temperature. As $L$ increases, the peak in $\delta O$ sharpens and approaches the critical temperature $T_{c}$ previously determined by large-scale MC simulations [36, 37]. Given the limited system sizes accessible in this study—and because a precise determination of critical properties of these models is beyond our present scope—we did not attempt a data collapse to extract $T_{c}$ or the critical exponents $\nu$ and $\beta$.
Appendix D Continuous spin models
@@
In this appendix, we study two paradigmatic models with continuous spin variables on the square lattice: the celebrated XY model and the classical Heisenberg model. The XY model undergoes a Berezinskii-Kosterlitz-Thouless (BKT) phase transition [39, *Kosterlitz_1973, *Kosterlitz_1974] at a critical temperature $T_{c}\approx 0.894$ [38], separating a low-temperature phase with quasi-long-range order from a high-temperature disordered phase characterized by an exponential decay of correlations. In contrast, for the 2D classical Heisenberg model, the existence of a BKT-like transition at finite temperature has been suggested but remains a subject of ongoing debate [42, 43, 44].
Figures 13 and 14 present preliminary MC results for the XY and classical Heisenberg model on the square lattice, where we compute the temperature dependence of the order parameter $O$ and its fluctuations $\delta O$ for system size up to $L=64$. A notable feature in the behavior of $O$ is observed: unlike previous cases involving systems with fully or partial orders at low temperatures—where $O$ gradually saturates to a finite value and exhibits a clear plateau—here, although $O$ approaches to unity as $T\to 0$, it does so with a significant slope. We interpret this as an indication of the absence of long-range order at any nonzero temperature, consistent with the celebrated Mermin–Wagner–Hohenberg theorem [45, *Hohenberg_1967]. Moreover, the XY and Heisenberg models exhibit remarkable similar behavior. We note that our MC results for $O$ and $\delta O$ closely resemble those previously reported for the magnetization $M$ and susceptibility $\chi$ [42]. It would be worthwhile to pursue more systematic studies at larger system sizes in the future to further substantiate these observations.